OpenStudy (anonymous):

if a circle has a diameter of of 16 cm , and i need to find the ARC length of a sector within the circle that is 135 degrees , how do i figure out the arc length?

OpenStudy (anonymous):



OpenStudy (campbell_st):

the fraction of the circle is based on the angle subtended is \[\frac{135}{360} = \frac{3}{8}\] circumference of a circle is \[C =\pi d\] the arc length is found by \[l = \frac{3}{8} \times \pi d\] just substitute your diameter measurement and evaluate

OpenStudy (ujjwal):

Find the angle in radians.. Now, angle in radians=(length of arc)/radius

OpenStudy (campbell_st):

you don't need to convert to radians... its the angle subtended over 360

OpenStudy (ujjwal):

Or follow unitary method: For 360 degrees, arc length= pi*d (i.e circumference) For 135 degrees, arc length= pi*d*135/360

OpenStudy (anonymous):

answer cold be 8pi correct they want me to answer a number and pi i gues how to get the answer

OpenStudy (anonymous):

could*

OpenStudy (anonymous):

6pi?

OpenStudy (anonymous):

9.42

OpenStudy (campbell_st):

thats if you use pi as 3.14

OpenStudy (campbell_st):

so its 3/8 x pi x 16 = 6pi cm which is correct or 6 x 3.14 = 18.84cm

OpenStudy (anonymous):

18.84 is why i wrote 8pi

OpenStudy (campbell_st):

well doing the calculation only gives 6 pi

OpenStudy (anonymous):

ok thx

OpenStudy (anonymous):

if the diamter was 8 it be 2pi

OpenStudy (campbell_st):

if the diameter was 8 the arc length is 3/8 x pi x 8 = 3pi

OpenStudy (anonymous):

The total circumference is 2*pi*r = 50.3 Now for 135 deg arc, it would be 135*50.3/360 That is 18.86

OpenStudy (anonymous):

If the diameter of the circle is 8 cm, then the radius is 4 cm. its 2 pi .

OpenStudy (campbell_st):

ok... 2 formula for the circumference of a circle option 1 using d = diameter \[C = \pi \times d\] Option 2 using r = radius \[C = 2\times \pi \times r\] and given d = 2r or the diameter = 2 radii I used the 1st formula... as the diameter was the given measurement.

OpenStudy (campbell_st):

here is a simple formula for you to follow in calculating the arc length l \[l = \frac{\theta}{360} \times Circumference\]

OpenStudy (anonymous):

sorry thanks

OpenStudy (campbell_st):

and theta is the angle at the centre

OpenStudy (campbell_st):

thats ok... just hope it helped